Mathematics 134 – Intensive Calculus for Science 2 Discussion 3 – Some interesting integrals! Wednesday January 31, 2001

Background

Today, we want to practice using the substitution method to evaluate some indefinite integrals. Instead of writing up a set of solutions to be handed in, each group will put their work on the blackboard, and we will discuss the answers as a group. Group 3 A) Use the substitution u = x^{2} + 1 to evaluate

∫ 4x(x^{2} + 1) dx

B) Evaluate

∫ sin(2 − 5t) dt

by an appropriate substitution. C) Use the substitution u = cos(x) to evaluate

∫ sin(x) cos(x) dx